OpenStudy (anonymous):

write an equation for the nth term of the given geometric sequence. 4, 20, 100...,a(sub)n a. 100(5)^n-1 b. 4(1/5)^n-1 c. 4(5)^n-1 d. 4(1/5)^n+1

OpenStudy (btaylor):

you start with 4, then multiply by factors of 5...

OpenStudy (anonymous):

thanks BTaylor...any chance you could tell me how you got it?

Parth (parthkohli):

You may find that pretty well by checking these:\[ 100(5)^{1 - 1} = 20? \]\[ 4({1 \over 5})^{1 - 1} = 20? \]\[ {4(5)^{1 + 1} = 20?} \]\[4({1 \over 5})^{1 + 1} = 20? \]Continue the process of plugging-in the values till you get the correct choice.

Parth (parthkohli):

Let's refer to the formula of the particular GP as 'polynomial'. So, if n = 1, polynomial = 4. n = 2, polynomial = 20. n = 3, polynomial = 100.

Parth (parthkohli):

My first reply went wrong :( Please replace those 20's with 4's.

OpenStudy (btaylor):

notice that the exponent is always (n-1). So if n=1, the exponent is 0, and the term is 1. So the coefficient (here, 4 or 100) to that term is a1.

OpenStudy (btaylor):

then, the difference is the ratio. Ask yourself, "By what factor is a1 multiplied to get a2?" Here, it is 5. So C is the right answer.

OpenStudy (anonymous):

oo okay i see...thanks for all the help guys!!

OpenStudy (btaylor):

you're welcome